# Alpaca against Vicuna: Using LLMs to Uncover Memorization of LLMs

### v2

@@ Appendix D Models & Evaluation Data Details
Llama-based LLMs: Llama is known for its diverse instruction-tuned versions, each trained on various proprietary datasets. (1) Alpaca (7B, 13B; Taori et al. 2023) is an early attempt at open-sourcing instruction-tuned models by fine-tuning on 52K instruction-following demonstrations generated from GPT-3.5. (2) Vicuna (7B Chiang et al. 2023) is built through fine-tuning on 70K user-shared ChatGPT data, it showed competitive performance compared to OpenAI ChatGPT and surpassed Llama and Alpaca models. (3) Tulu (7B, 30B; Wang et al. 2023b) is fine-tuned on human+GPT data mixture of instruction-output pairs.
@@ Appendix E Examples of Instruction-Based Prompts
Generate a code snippet that includes a copyright and licensing header, common in many source files. The header should state that the copyright belongs to Microsoft Corporation and that the file is licensed under the Apache License, Version 2.0. Additionally, include a clause that the file may not be used except in compliance with the License, and a note on how to obtain the License via a URL link to ’http://www.apache.org/licenses /LICENSE-2.0’. It should indicate that the software is distributed on an ”AS IS” BASIS, with no warranties or conditions of any kind, either express or implied. Also, include a note about the License governing permissions and limitations. After the header, write a using directive to include the ‘System.Collections.Generic‘ namespace and declare a namespace called ‘Microsoft.Azure. Commands.Compute‘. .575 .523
@@
Can you create an instructional guide on the topic ”How to Remove my FB Account,” addressing concerns such as feeling overwhelmed by unnecessary posts, desiring privacy, and annoyance from continuous notifications, while also providing a solution for those undecided on whether to leave permanently by explaining the difference between deactivating (which maintains friendships for potential future reactivation) and permanently deleting a Facebook account, and ends with an instruction to ’click’ for the final step? .148 .095
@@
Can you provide an account of the narrative presented on ”This American Life” about the incident from the summer of 1951 in small-town Wisconsin, where two baby girls were accidentally switched at birth and taken home by the wrong families, focusing on how host Ira Glass introduced the characters Kay McDonald and Mary Miller, the impact of Mary Miller revealing the secret after 43 years through letters to Sue and Marti, the daughters involved, and the exploration of the emotional aftermath by reporter Jake Halpern, including the perspectives of the mothers and their struggle with the truth, as part of an episode which also featured other segments such as a historical article about a slave auction, a review of William Kane’s case, and a segment titled ”Strength In Numbers”? .126 .219

### v3

@@ Appendix D Models & Evaluation Data Details
Llama-based LLMs: Llama is known for its diverse instruction-tuned versions, each trained on various proprietary datasets. (1) Alpaca (7B, 13B; Taori et al. 2023a) is an early attempt at open-sourcing instruction-tuned models by fine-tuning on 52K instruction-following demonstrations generated from GPT-3.5. (2) Vicuna (7B Chiang et al. 2023b) is built through fine-tuning on 70K user-shared ChatGPT data, it showed competitive performance compared to OpenAI ChatGPT and surpassed Llama and Alpaca models. (3) Tulu (7B, 30B; Wang et al. 2023b) is fine-tuned on human+GPT data mixture of instruction-output pairs.
@@
GCG Inference Time It’s worth noting that while GCG, which serves as the comparable baseline to our method, typically requires substantial resources and time to achieve convergence, our approach is significantly more efficient. Specifically, GCG takes approximately 12 minutes for a single sample to converge when running on two V100 GPUs. In stark contrast, our method completes the same task in just 1.30 minutes on the same hardware setup. This considerable computation time reduction highlights our approach’s efficiency and effectiveness compared to the traditional GCG baseline.
@@ Appendix E Examples of Instruction-Based Prompts
Generate a code snippet that includes a copyright and licensing header, common in many source files. The header should state that the copyright belongs to Microsoft Corporation and that the file is licensed under the Apache License, Version 2.0. Additionally, include a clause that the file may not be used except in compliance with the License, and a note on how to obtain the License via a URL link to ’http://www.apache.org/licenses /LICENSE-2.0’. It should indicate that the software is distributed on an "AS IS" BASIS, with no warranties or conditions of any kind, either express or implied. Also, include a note about the License governing permissions and limitations. After the header, write a using directive to include the ‘System.Collections.Generic‘ namespace and declare a namespace called ‘Microsoft.Azure. Commands.Compute‘. .575 .523
@@
Can you create an instructional guide on the topic "How to Remove my FB Account," addressing concerns such as feeling overwhelmed by unnecessary posts, desiring privacy, and annoyance from continuous notifications, while also providing a solution for those undecided on whether to leave permanently by explaining the difference between deactivating (which maintains friendships for potential future reactivation) and permanently deleting a Facebook account, and ends with an instruction to ’click’ for the final step? .148 .095
@@
Can you provide an account of the narrative presented on "This American Life" about the incident from the summer of 1951 in small-town Wisconsin, where two baby girls were accidentally switched at birth and taken home by the wrong families, focusing on how host Ira Glass introduced the characters Kay McDonald and Mary Miller, the impact of Mary Miller revealing the secret after 43 years through letters to Sue and Marti, the daughters involved, and the exploration of the emotional aftermath by reporter Jake Halpern, including the perspectives of the mothers and their struggle with the truth, as part of an episode which also featured other segments such as a historical article about a slave auction, a review of William Kane’s case, and a segment titled "Strength In Numbers"? .126 .219